if you have $30.00 for one year at 1.5 intrest rate how much will you have...
Answer by stanbon(75887)

if you have $30.00 for one year at 1.5 interest rate how much will you have?
-----
You will have 30 + 0.015*30 = 30 + 0.45 = $30.45
===================================================
Cheers,
Stan H.